OpenStudy (aravindg):

1.Water drops from the roof of a building 10 m high.The drops fall at regular intervals of time, the first drop striking the floor at the instant the 4th drop begins to fall.Find positions of individual drops,when a drop strikes the floor. 2.A Launch travels across a river from point A to a point B on the opposite bank along the straight line AB forming an angle α with the bank.The flag on the mast of launch makes an angle β with its direction of motion.Determine speed of launch with respect to bank.The velocity of wind is u perpendicular to the stream. 3. Two intersecting straight lines move

OpenStudy (anonymous):

the first drop will be at h=0m the second at h=3.3m the third at h=6.6m and the forth at h=10m. no clue abut the second uestion though. too much trig form me. may come back though if i have an idea

OpenStudy (anonymous):

The time intervals will be the same not the heights dropped. Time taken for the first drop to hit the ground can be found from: s=ut +1/2 at^2 10=0 + 1/2 . g . t^2 Using 10 for gravity, or 9.81 if you are being that accurate... 10 = 1/2 . 10. t^2 So t = sqrt(2) or 1.4 seconds. Does that sound right? Yes. So the first drop will hit the ground after sqrt(2) seconds. The second drop will have dripped 2/3 of that time before; the third 1/3 of that time; and the fourth will just be starting. Plug these times back into new s=ut +1/2 at^2 and you will get individual "s" for each drop. s1=10; s2=40/9; s3=10/9; s4=0.